c) A capillary tube with an inside radius of 638 nm can support a 100 mm column of liquid that has a density of 820 kg * m ^ - 3 . The observed contact angle is 12.5 degrees . Find the surface tension of the liquid for h = 20 cm .

1
Expert's answer
2021-12-07T09:56:16-0500

According to the Jurin's law


"h=\frac{2\sigma}{(\rho-\rho_0)gr}\cdot\cos\theta\to \sigma=h(\rho-\rho_0)gr/(2\cos\theta)="


"=0.2\cdot(820-1.29)\cdot9.8\cdot638\cdot10^{-9}/(2\cdot\cos12.5°)=5.24\cdot10^{-4} \ (N/m)" . Answer